### v5

# Breaking the Exploration Bottleneck: Rubric-Scaffolded Reinforcement Learning for General LLM Reasoning

Abstract
Recent advances in Large Language Models (LLMs) have underscored the potential of Reinforcement Learning (RL) to facilitate the emergence of reasoning capabilities. Despite the encouraging results, a fundamental dilemma persists as RL improvement relies on learning from high-quality samples, yet the exploration for such samples remains bounded by the inherent limitations of LLMs. This, in effect, creates an undesirable cycle in which what cannot be explored cannot be learned. In this work, we propose Rubric-Scaffolded Reinforcement Learning (RuscaRL), a novel instructional scaffolding framework designed to break the exploration bottleneck for general LLM reasoning. Specifically, RuscaRL introduces checklist-style rubrics as (1) explicit scaffolding for exploration during rollout generation, where different rubrics are provided as external guidance within task instructions to steer diverse high-quality responses. This guidance is gradually decayed over time, encouraging the model to internalize the underlying reasoning patterns; (2) verifiable rewards for exploitation during model training, where we can obtain robust LLM-as-a-Judge scores using rubrics as references, enabling effective RL on general reasoning tasks. Extensive experiments demonstrate the superiority of the proposed RuscaRL across various benchmarks, effectively expanding reasoning boundaries under the Best-of-N evaluation. Notably, RuscaRL significantly boosts Qwen2.5-7B-Instruct from 23.6 to 50.3 on HealthBench-500, surpassing GPT-4.1. Furthermore, our fine-tuned variant on Qwen3-30B-A3B-Instruct achieves 61.1 on HealthBench-500, outperforming leading LLMs including OpenAI-o3. Our code is available at https://github.com/IANNXANG/RuscaRL.
##footnotetext: Corresponding author.
1 Introduction
Large Language Models (LLMs) have demonstrated tremendous potential over a wide spectrum of complex reasoning tasks, including legal analysis (Choi et al., 2021; Lai et al., 2024; Fei et al., 2023; Trautmann et al., 2022), robotic manipulation (Driess et al., 2023; Zitkovich et al., 2023; Firoozi et al., 2025; Zhou et al., 2023a), and software development (Anysphere, 2023; Fan et al., 2023; Hou et al., 2024). However, advancing the general reasoning capabilities of LLMs remains a significant challenge (Zhao et al., 2023; Huang & Chang, 2022). To address this, recent breakthroughs in Reinforcement Learning with Verifiable Rewards (RLVR), as exemplified by DeepSeek-R1 (Guo et al., 2025) and OpenAI-o3 (OpenAI, 2025a), have proven that leveraging verifiable rewards as feedback signals can successfully facilitate the emergence of sophisticated reasoning capabilities in LLMs (Lambert et al., 2024; Yang et al., 2025; Kimi et al., 2025).
Despite the encouraging results, conventional RLVR tends to be more applicable to domains with objectively verifiable answers. For instance, in areas such as mathematical proof (Ren et al., 2025; Chen et al., 2025) and code generation (Qwen, 2025; Le et al., 2022), correctness can be explicitly determined through formal proof verification or automated unit tests. In these contexts, the reward signal is clear and well-aligned with the task objective, allowing RLVR to effectively guide models toward correct solutions. Unfortunately, many real-world tasks like medical consultation (Lin et al., 2025; Singhal et al., 2023; Zhang et al., 2023) and creative writing (Wu et al., 2025c; Franceschelli & Musolesi, 2024) are inherently open-ended. These tasks often require multidimensional evaluation and lack a single, verifiable ground-truth answer. To tackle this problem, several very recent concurrent works (Arora et al., 2025; Kimi et al., 2025; Gunjal et al., 2025; Viswanathan et al., 2025; Huang et al., 2025; Dou et al., 2025) have explored rubric-based evaluation that decomposes desirable responses into checklist-style criteria (e.g., factuality, coherence, completeness). By leveraging LLM-as-a-Judge to score each criterion and aggregating results into scalar rewards, rubrics provide more stable and reliable feedback signals suitable for RLVR in open-ended domains.
Nevertheless, as shown in the left side of Figure 1, a fundamental exploration bottleneck remains as RL requires high-quality samples to improve, yet exploration for such samples remains bounded by the inherent capabilities of LLMs (Yue et al., 2025; Wu et al., 2025a; Liu et al., 2025b; Dong et al., 2025). This creates an inevitable loop where the inability to explore restricts the ability to learn. A growing line of work has attempted to enhance exploration in RLVR for LLMs (Liu et al., 2025b; a; Dong et al., 2025; Zheng et al., 2025; Lei et al., 2025; Li et al., 2025; Cheng et al., 2025). However, these methods largely bias the policy distribution toward high-reward responses already supported by base models, rather than truly expanding its reasoning boundaries (Wu et al., 2025a). Worse still, RL itself has a natural tendency to narrow the exploration space: policy entropy gradually collapses during training, causing the model to converge toward a limited set of reasoning trajectories (Zhao et al., 2025; Yue et al., 2025; Wu et al., 2025a; Yu et al., 2025; Liu et al., 2025b). This, in turn, undermines the potential of RLVR to explore more diverse and higher-quality solutions.
Figure 1: (Left) A conceptual illustration of exploration bottleneck and scaffolding. (Right) Performance comparison of different LLMs on HealthBench-500. Note that Qwen3-30B-Instruct denotes Qwen3-30B-A3B-Instruct, and Qwen3-235B-Thinking denotes Qwen3-235B-A22B-Thinking.
In this work, we introduce Rubric-Scaffolded Reinforcement Learning, termed as RuscaRL, which employs a novel instructional scaffolding framework to break the exploration bottleneck of RLVR. Technically, RuscaRL leverages rubrics in two complementary ways: (1) Explicit scaffolding during rollout generation. For each instruction, RuscaRL generates a group of candidate responses by using rubrics as external guidance. We propose intra-group scaffolding differentiation to provide varying levels of rubrics within each group, enabling diverse and high-quality responses. To further internalize underlying reasoning patterns, we use inter-step scaffolding decay to gradually remove the scaffolding over training, thereby minimizing reliance on external guidance. (2) verifiable rewards during model training. The model responses are assessed based on multiple criteria derived from rubrics. For each criterion, we employ a Grader LLM that performs a binary evaluation (i.e., true or false), determining whether the response satisfies that specific requirement. The outcomes are then combined through aggregation to obtain a robust reward signal, facilitating effective RL across different general tasks. Our main contributions are summarized as follows:
  • We introduce instructional scaffolding as a novel paradigm in RLVR for LLMs, which pioneers the integration of external guidance within task instructions to improve rollout diversity and quality, thereby enabling more efficient exploration during RL.
  • We propose Rubric-Scaffolded Reinforcement Learning (RuscaRL), an innovative RLVR framework designed to break the exploration bottleneck, integrating checklist-style rubrics as both explicit scaffolding for exploration and verifiable rewards for exploitation.
  • Extensive experiments demonstrate that RuscaRL yields results superior to state-of-the-art counterparts. Notably, RuscaRL enables small LLMs (e.g., Qwen3-30B) to achieve performance on par with leading LLMs (e.g., OpenAI-o3) on HealthBench-500, as shown in the right side of Figure 1.
2 Related Works
Rubric-based Methods. Rubrics are structured evaluation frameworks that decompose complex assessment tasks into specific, verifiable criteria. To address general task evaluation, rubric-based evaluation approaches have emerged across medical (Arora et al., 2025), code (Pathak et al., 2025), and other domains (Fan et al., 2025; Galvan-Sosa et al., 2025; Winata et al., 2025). Building upon these frameworks, researchers apply rubrics as reward signals in RL (Kimi et al., 2025; Gunjal et al., 2025; Viswanathan et al., 2025), using LLMs as graders to provide fine-grained feedback for tasks lacking ground truth. This approach has shown promising results across LLM alignment (Dineen et al., 2025), instruction following (Viswanathan et al., 2025), and open-ended QA (Gunjal et al., 2025; Huang et al., 2025; Dou et al., 2025; Kimi et al., 2025).
Exploration in RL for LLMs. Existing RL methods face insufficient exploration in complex reasoning tasks, with policies trapped in local optima and reasoning boundaries collapsing (Wu et al., 2025a; Yue et al., 2025; Liu et al., 2025b). Current solutions include prolonged training (Liu et al., 2025b; a), entropy-based exploration (Dong et al., 2025; Zheng et al., 2025; Lei et al., 2025; Li et al., 2025; Cheng et al., 2025), and external guidance (Zhang et al., 2025a; RRV et al., 2025), but these approaches fail to break the exploration bottleneck because they either explore within the initial policy distribution or provide only coarse directional signals without structured intermediate guidance. In contrast, RuscaRL provides explicit scaffolding via checklist-style rubrics, guiding trajectories with verifiable criteria while enabling scaffolding decay for pattern internalization.
3 Preliminary
RL Algorithms for LLMs. We adopt Group Relative Policy Optimization  (GRPO) (Shao et al., 2024) as our core RL algorithm for training language models with rubric-based rewards. Unlike Proximal Policy Optimization (PPO) (Schulman et al., 2017), GRPO eliminates the need for a value model by using group-based advantage estimation. For each instruction $q\sim\mathcal{D}$, where $\mathcal{D}$ denotes the distribution over the training dataset $\mathcal{D}$, GRPO samples a group of $G$ responses $\{o_{1},o_{2},\ldots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and optimizes the policy $\pi_{\theta}$ by maximizing the following objective:
$\begin{split}\mathcal{J}_{\text{GRPO}}\left(\theta\right)&=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|q)}\\ &\quad\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\left(\rho_{i,t}(\theta)\hat{A}_{i},\text{clip}\left(\rho_{i,t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right)\right],\end{split}$ (1)
where $o_{i}$ is a response sampled from the old policy $\pi_{\theta_{\text{old}}}$ given instruction $q$, $t$ denotes the token position within response $o_{i}$, $\rho_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ is the token-level importance ratio between the current policy and the previous policy, and $\epsilon$ is the clipping coefficient (Schulman et al., 2017). The group-relative advantage is computed as:
$\hat{A}_{i}=\frac{r_{i}-\operatorname{mean}\big(\left\{r_{j}\right\}_{j=1}^{G}\big)}{\operatorname{std}\big(\left\{r_{j}\right\}_{j=1}^{G}\big)},$ (2)
where $r_{i}$ is the reward for response $o_{i}$, and the advantage is normalized using the mean and standard deviation of the $G$ sampled rewards.
4 Methodology
Figure 2: Overview of the RuscaRL framework: (a) Rubric data sample with criteria and points, (b) Rubric-scaffolded reinforcement learning with reward computation and policy updates, (c) Intra-group scaffolding differentiation and inter-step scaffolding decay mechanisms.
To address the exploration bottleneck problem, we propose the RuscaRL framework, as illustrated in Figure 2. RuscaRL leverages rubrics in two complementary ways: (1) Explicit scaffolding during rollout generation, where the model generates candidate responses using rubrics as external guidance with intra-group differentiation and inter-step decay (Section 4.2); (2) Verifiable rewards during model training, where responses are assessed based on multiple criteria derived from rubrics through binary evaluation and aggregation (Section 4.3). In the following, we first introduce the basic concept of rubrics, then detail these two core components.
4.1 Rubric-based Evaluation System
Evaluating complex and open-ended tasks is inherently challenging, as responses often differ in structure, style, and content, making it difficult for rule-based evaluation to provide reliable assessments. To address this gap, recent work (Arora et al., 2025) has proposed rubric-based evaluation, which specifies fine-grained, multi-dimensional criteria that can be applied at scale. This design combines the objectivity of automatic metrics with the principled guidance of structured standards, yielding robust scores that better capture response quality, coherence, and completeness.
Formally, a rubric $\mathcal{R}=\{c_{1},c_{2},\ldots,c_{N}\}$ is defined as a set of $N$ verifiable criteria. Each criterion $c_{i}$ is specified by a clear description and corresponding points $p_{i}$ indicating its contribution to the overall evaluation. We define the point vector as $\mathbf{p}=[p_{1},p_{2},\ldots,p_{N}]$. For example, given the instruction “Can you introduce the idea of reinforcement learning?”, criteria may include: “step-by-step analysis with complete structural logic” (+3 points), “explain key elements: agent, environment, reward” (+6 points), and negative items like “confuse environment and reward roles in explanation” (-7 points). Points are added or subtracted depending on whether each criterion is satisfied.
Given an instruction $q$ and its corresponding rubric $\mathcal{R}$ both sampled from the data distribution $\mathcal{D}$, and a model response $o$ generated through the policy model $\pi_{\theta}(o|q)$, we first construct a judge prompt for each criterion $c_{i}$ by combining the instruction $q$, response $o$, and criterion $c_{i}$. The judge prompt template for the grader is provided in Appendix E.1. For a single criterion evaluation, the grader function $\mathcal{G}$ implemented by a LLM (Zheng et al., 2023b; Gu et al., 2024) takes the judge prompt as input and outputs a binary decision $b_{i}=\mathcal{G}(q,o,c_{i})\in\{0,1\}$ indicating whether criterion $c_{i}$ is satisfied (true or false). Extending this to the full rubric, the grader evaluates all criteria and produces a binary indicator vector $\mathbf{b}=\mathcal{G}(q,o,\mathcal{R})=[b_{1},b_{2},\ldots,b_{N}]$ where each $b_{i}$ represents the satisfaction of criterion $c_{i}$. The final score vector is obtained by element-wise multiplication: $\mathbf{s}=\mathbf{b}\odot\mathbf{p}=[b_{1}p_{1},b_{2}p_{2},\ldots,b_{N}p_{N}]$, providing fine-grained score across all specified criteria. We also compute the total possible score $S_{total}=\sum_{j=1}^{M}p_{j}$ where $M$ is the number of positive-point criteria, which will be used for normalization in the reward calculation.
4.2 Rubric-based Scaffolding Mechanism for RL Exploration
During RL training on complex reasoning tasks, models often fail to sustain effective exploration (Wu et al., 2025a; Yue et al., 2025; Liu et al., 2025b): initial randomness quickly diminishes, policy entropy collapses, and the model prematurely converges to suboptimal reasoning patterns. This collapse severely limits the discovery of diverse and high-quality solution trajectories.
To mitigate this issue, we draw inspiration from instructional scaffolding theory in educational psychology (Vygotsky & Cole, 1978). According to Vygotsky’s Zone of Proximal Development, when learners’ capabilities are insufficient, they benefit from structured guidance to bridge the gap between current ability and target performance. As competence grows, this scaffolding should be gradually withdrawn to foster independent problem-solving (Wood et al., 1976).
Building on this insight, we design a rubric-based scaffolding mechanism that provides varying numbers of rubric criteria as explicit guidance throughout the training process, helping models gradually learn to generate high-quality responses. Specifically, our rubric-based scaffolding mechanism augments the original policy function by adding a subset of rubric criteria $\mathcal{R}_{S}$ as additional guidance, representing the policy as $\pi_{\theta}(o|q,\mathcal{R}_{S})$. The specific prompt template for incorporating scaffolding is detailed in Appendix E.2. Additionally, we design a two-dimensional control mechanism to determine the rubrics scaffolding ratio $\lambda_{S}$, and then sample criteria from the complete rubric set $\mathcal{R}$ to form $\mathcal{R}_{S}$, i.e., $|\mathcal{R}_{S}|=\text{round}(\lambda_{S}\times|\mathcal{R}|)$. We instantiate this mechanism in two dimensions: intra-group Scaffolding differentiation and inter-step Scaffolding decay.
Intra-Group Scaffolding Differentiation. In RL algorithms with multi-sampling, such as GRPO, computing group-relative advantages (Eq. 2) requires response diversity to avoid collapse into homogeneous samples. To this end, we assign different levels of rubric scaffolding within each group, encouraging both guided and independent exploration. Concretely, we define a group-level ratio vector $\bm{\lambda}_{group}=[\lambda_{1},\lambda_{2},\ldots,\lambda_{G}]$ where $\lambda_{i}=\frac{G-i}{G-1}$ for the $i$-th sample in the group of size $G$. This linear differentiation ensures that some samples benefit from stronger scaffolding while others are deliberately exposed to weaker guidance, thereby enhancing intra-group diversity.
Inter-Step Scaffolding Decay. Inspired by instructional scaffolding theory, we gradually reduce guidance as the model develops independent learning strategies using a sigmoid function $\lambda_{step}(t)=\frac{1}{1+e^{\alpha(t-t_{0})}}$, where $t$ is the current training progress ($t\in[0,1]$), $t_{0}$ is the midpoint, and $\alpha$ controls the steepness of decay. This mechanism prevents overfitting to external guidance by creating an adaptive learning environment where the model initially benefits from guidance to overcome the exploration bottleneck, then gradually transitions to independent reasoning as capabilities mature.
Integrated Scaffolding Mechanism. Finally, we combine intra-group differentiation and inter-step decay into a unified ratio vector:
$\bm{\lambda}_{S}=\lambda_{step}\left(t\right)\times\bm{\lambda}_{group}=\left[\lambda_{S,1},\lambda_{S,2},\ldots,\lambda_{S,G}\right],$ (3)
where $\lambda_{S,i}=\lambda_{step}(t)\times\lambda_{i}=\frac{1}{1+e^{\alpha(t-t_{0})}}\times\frac{G-i}{G-1}$ represents the scaffolding ratio for the $i$-th sample in the group. This integrated mechanism simultaneously promotes response diversity within each group and adaptively reduces reliance on scaffolding across training steps, jointly addressing the problems of homogeneity and overfitting.
4.3 Rubric-based Reward Function for RL Exploitation
To provide robust and reliable reward signals for general reasoning tasks, we design rubric-based reward functions. The multi-dimensional score vector $\mathbf{s}=[s_{1},s_{2},\ldots,s_{N}]$ obtained from the evaluation system is aggregated into a final scalar reward by directly summing all criterion scores and normalizing by the total possible score computed in Section 4.1:
$S=\frac{\sum_{i=1}^{N}s_{i}}{S_{total}},$ (4)
where $S$ represents the final score, $s_{i}$ is the score of the $i$-th criterion, and $S_{total}$ is the total possible score of all positive-point criteria computed in Section 4.1. This calculation method results in scores that fall within the interval $[0,1]$ in most cases, with occasional negative scores possible. We directly adopt this rubric-based score $S$ as our reward: $r_{i}=S_{i}$, where $r_{i}$ is the reward for the $i$-th response. This approach enables application to open-ended tasks without ground truth answers while providing more robust assessment than holistic LLM scoring.
Once the rubric-based rewards are obtained, we employ them to train the policy model using RL algorithms. The training process follows the policy gradient framework, where the model learns to maximize the expected reward. Algorithm 1 in Appendix B outlines the complete training procedure. Additionally, to help the model better internalize underlying reasoning patterns, the log probability computation in training is based on $\pi_{\theta}(o_{i,t}|q,o_{i,<t})$ rather than $\pi_{\theta}(o_{i,t}|q,\mathcal{R}_{S},o_{i,<t})$. For detailed analysis on importance sampling, see Appendix D.5.
5 Experiments
To demonstrate the effectiveness of the proposed RuscaRL method, we conduct experiments across multiple benchmarks spanning medical, writing, instruction following, and STEM domains. Our experiments seek to answer the following questions: (1) Does RuscaRL demonstrate consistent effectiveness across different models and tasks, and how does it compare against existing fine-tuning methods? (Section 5.2 and Appendix D.1, D.2, D.3) (2) How does RuscaRL break the exploration bottleneck in RL for LLM reasoning? (Section 5.3 and Appendix D.4) (3) What is the impact of different components in the rubric-based scaffolding mechanism? (Section 5.4 and Appendix D.5)
5.1 Experimental Setups
Table 1: Main results comparison with closed-source and open-source models across different benchmarks. The best results in each box are highlighted in bold.
Model Medical Writing Instruction Following
HealthBench-500 LLMEval-Med MedQA MedMCQA WritingBench Creative Writing IFEVAL IFBench
Closed-source and Open-source Models
OpenAI-o3 59.8 78.9 96.0 84.7 77.7 81.4 91.6 67.8
GPT-4.1 47.9 71.2 92.4 80.0 69.0 79.0 87.0 37.2
gpt-oss-20b 42.5 68.8 85.6 68.1 66.6 39.4 83.5 48.7
Rubicon-Preview 50.8 73.2 85.1 70.7 73.0 66.4 82.4 33.4
Our Models
Qwen3-30B-A3B-Instruct 46.8 71.4 84.2 71.3 78.4 74.4 83.0 31.9
+ RuscaRL 61.1 (+14.3) 73.0 (+1.6) 84.8 (+0.6) 71.9 (+0.6) 79.2 (+0.8) 74.3 (-0.1) 84.5 (+1.5) 32.1 (+0.2)
Qwen3-30B-A3B-Base 11.4 43.3 73.6 65.1 36.8 35.8 39.0 13.3
+ RuscaRL 48.3 (+36.9) 60.9 (+17.6) 71.3 (-2.3) 65.4 (+0.3) 59.0 (+22.2) 46.0 (+10.2) 76.3 (+37.3) 30.3 (+17.0)
Qwen2.5-7B-Instruct 23.6 47.9 61.8 56.3 45.3 37.4 71.0 26.8
+ RuscaRL 50.3 (+26.7) 61.2 (+13.3) 63.5 (+1.7) 56.5 (+0.2) 56.3 (+11.0) 38.6 (+1.2) 75.3 (+4.3) 31.0 (+4.2)
Qwen2.5-7B 8.2 28.2 55.3 55.0 23.5 30.3 32.0 14.5
+ RuscaRL 46.2 (+38.0) 47.8 (+19.6) 58.2 (+2.9) 55.6 (+0.6) 45.8 (+22.3) 34.8 (+4.5) 56.2 (+24.2) 25.9 (+11.4)
Llama-3.1-8B-Instruct 12.4 29.8 66.8 58.0 36.7 44.5 72.6 22.6
+ RuscaRL 46.0 (+33.6) 46.3 (+16.5) 70.7 (+3.9) 60.7 (+2.7) 52.7 (+16.0) 54.2 (+9.7) 79.7 (+7.1) 31.1 (+8.5)
Llama-3.1-8B 0 9.1 36.9 35.9 12.9 26.3 18.1 11.6
+ RuscaRL 25.8 (+25.8) 29.5 (+20.4) 49.7 (+12.8) 45.4 (+9.5) 35.6 (+22.7) 33.3 (+7.0) 55.6 (+37.5) 21.4 (+9.8)
Figure 3: STEM Benchmarks.
Models and Training Settings. We use multiple initial models from different series and parameter scales for our experiments, including both instruct models and base models from the Qwen2.5 series (Yang et al., 2024), the Qwen3 series (Yang et al., 2025), and the Llama-3 series (Meta-AI, 2025; Grattafiori et al., 2024). All models are trained using the GRPO algorithm on the verl framework (Sheng et al., 2025). Detailed training settings are provided in Appendix C.1.
Evaluation Benchmarks. We use HealthBench-500, a randomly selected subset of 500 samples from HealthBench (Arora et al., 2025), as our held-out evaluation set. Additionally, we evaluate on other medical benchmarks including LLMEval-Med (Zhang et al., 2025b), MedQA (Jin et al., 2021), and MedMCQA (Pal et al., 2022). For the Writing domain, we use WritingBench (Wu et al., 2025c) and Creative Writing v3 (Paech, 2025) benchmarks. For the Instruction Following domain, we use IFEVAL (Zhou et al., 2023b) and IFBench (Pyatkin et al., 2025) benchmarks. For the STEM domain, we use GPQA-Diamond (Rein et al., 2024), MMLU (Hendrycks et al., 2020), MMLU-Pro (Wang et al., 2024), MATH-500 (Lightman et al., 2023), AMC 2023111https://huggingface.co/datasets/knoveleng/AMC-23, AIME 2024, and AIME 2025222https://artofproblemsolving.com/wiki/index.php/AIME_Problems_and_Solutions. Detailed evaluation settings are provided in Appendix C.2.
Baselines. We compare RuscaRL against four representative baseline methods: (1) Rubric-based RL: A rubric-as-reward RL baseline implemented with GRPO algorithm using rubrics scores as rewards (Gunjal et al., 2025). (2) Rubric-based RL with full scaffolding (Rubric-based RL-S): A method that provides all rubrics with scaffolding support in the instruction, without intra‑group differentiation and without an inter‑step decay function. (3) SFT: Fine-tuned on GPT-4.1 (OpenAI, 2025b) demonstrations generated with scaffolding support. (4) SFT + Rubric-based RL: A combination approach that first applies SFT and then applies rubric-based RL training.
5.2 Overall Performance
RuscaRL achieves consistent and notable gains across tasks and model scales, showcasing its effectiveness and broad generalization. Across medical, writing, and instruction-following tasks (Table 1), RuscaRL achieves significant gains over multiple initial models, with Qwen3-30B-A3B-Instruct on HealthBench-500 surpassing many leading models (e.g., OpenAI-o3). Notably, RuscaRL is particularly effective on instruct models and provides larger benefits for weaker models, such as Llama-3.1-8B-Instruct. This advantage stems from our scaffolding approach, which leverages the model’s existing instruction-following ability to elicit higher-quality and more diverse responses, explaining why RuscaRL is especially well-suited for training on models with strong instruction-following capacities. Meanwhile, RuscaRL has also been successfully extended to the STEM domain: experiments on Qwen2.5-7B-Instruct show consistent performance improvements across all STEM benchmarks (see Figure 3). More detailed results about performance across different model series and scales are provided in Appendix D.1, which further demonstrates the robustness and broad applicability of our approach. Additionally, we explore the effects of mixing training data from different domains in Appendix D.2.
Table 2: Performance comparison of different baseline methods and RuscaRL across multiple domains. Bold indicates the best performance of each method.
Method Medical Writing Instruction Following
HealthBench-500 LLMEval-Med MedQA MedMCQA WritingBench Creative Writing IFEVAL IFBench
Qwen2.5-7B-Instruct
Initial 23.6 47.9 61.8 56.3 45.3 37.4 71.0 26.8
Rubric-based RL 41.2 (+17.6) 54.6 (+6.7) 62.1 (+0.3) 56.3 (+0.0) 53.7 (+8.4) 38.8 (+1.4) 75.1 (+4.1) 29.3 (+2.5)
Rubric-based RL-S 36.6 (+13.0) 56.1 (+8.2) 57.9 (-3.9) 52.4 (-3.9) 45.8 (+0.5) 38.3 (+0.9) 71.9 (+0.9) 28.6 (+1.8)
RuscaRL (Ours) 50.3 (+26.7) 61.2 (+13.3) 63.5 (+1.7) 56.5 (+0.2) 56.3 (+11.0) 38.6 (+1.2) 75.3 (+4.3) 31.0 (+4.2)
SFT 38.3 (+14.7) 52.6 (+4.7) 60.8 (-1.0) 57.3 (+1.0) 62.8 (+17.5) 45.3 (+7.9) 75.2 (+4.2) 25.2 (-1.6)
SFT + Rubric-based RL 55.1 (+31.5) 58.5 (+10.6) 59.7 (-2.1) 56.4 (+0.1) 66.7 (+21.4) 43.6 (+6.2) 82.1 (+11.1) 29.6 (+2.8)
SFT + RuscaRL (Ours) 56.9 (+33.3) 58.8 (+10.9) 61.6 (-0.2) 56.9 (+0.6) 66.8 (+21.5) 43.9 (+6.5) 82.5 (+11.5) 30.6 (+3.8)
Qwen2.5-7B
Initial 8.2 28.2 55.3 55.0 23.5 30.3 32.0 14.5
Rubric-based RL 41.9 (+33.7) 46.5 (+18.3) 48.2 (-7.1) 49.9 (-5.1) 40.2 (+16.7) 33.8 (+3.5) 53.4 (+21.4) 25.5 (+11.0)
Rubric-based RL-S 21.7 (+13.5) 44.4 (+16.2) 60.3 (+5.0) 55.5 (+0.5) 43.9 (+20.4) 25.7 (-4.6) 52.3 (+20.3) 20.4 (+5.9)
RuscaRL (Ours) 46.2 (+38.0) 47.8 (+19.6) 58.2 (+2.9) 55.6 (+0.6) 45.8 (+22.3) 34.8 (+4.5) 56.2 (+24.2) 25.9 (+11.4)
SFT 32.2 (+24.0) 40.0 (+11.8) 56.5 (+1.2) 54.4 (-0.6) 56.6 (+33.1) 42.5 (+12.2) 69.7 (+37.7) 20.8 (+6.3)
SFT + Rubric-based RL 36.4 (+28.2) 39.7 (+11.5) 57.1 (+1.8) 54.1 (-0.9) 57.1 (+33.6) 43.2 (+12.9) 71.6 (+39.6) 23.7 (+9.2)
SFT + RuscaRL (Ours) 35.4 (+27.2) 42.7 (+14.5) 58.2 (+2.9) 55.1 (+0.1) 57.4 (+33.9) 42.6 (+12.3) 72.0 (+40.0) 23.1 (+8.6)
RuscaRL consistently outperforms Rubrics-based methods across tasks. As shown in Table 2, in the direct RL setting, RuscaRL achieves the best performance on most medical, writing, and instruction-following tasks, delivering larger and more stable gains than Rubric-based RL and RL-S (e.g., 50.3 vs. 41.2 and 36.6 accuracy on HealthBench-500 with Qwen2.5-7B-Instruct). In the SFT-then-RL setting, both RuscaRL and Rubric-based RL achieve additional improvements over SFT, with RuscaRL generally yielding larger gains across most tasks, though the magnitude is smaller than in the direct RL setting. We argue that RuscaRL essentially leverages rubrics as prior knowledge to guide exploration, while SFT also serves to accelerate RL exploration (cold start). Since both mechanisms overlap in facilitating exploration, this may explain why the performance gap between RuscaRL and Rubric-based RL narrows under the SFT-then-RL setting.
5.3 Analysis
In this subsection, we conduct an analysis of RuscaRL using Qwen2.5-7B-Instruct as the initial model and HealthBench as the training and evaluation dataset. In addition, we compare three approaches in the following analysis: RuscaRL, RuscaRL (RuscaRL without the inter-step decay mechanism), and Rubric-based RL. We use the Best-of-N [^3] For cost considerations in the Best-of-N evaluation, we use Qwen3-32B (non-thinking) as the Grader LLM. metric to reflect both the model’s reasoning boundaries (at large N) and sampling efficiency (at small N).
Figure 4: Best-of-N Performance.
RuscaRL significantly improves sampling efficiency and reasoning boundaries. As shown in Figure 4, RuscaRL significantly improves single-sample quality at N=1, indicating that the scaffolding mechanism effectively enhances the model’s reasoning stability. At N=2048, its performance ceiling surpasses both the initial model and Rubric-based RL, validating its advantage in expanding the reasoning boundary. Moreover, RuscaRL exhibits a steeper performance curve across N, meaning it can achieve the same performance with fewer samples. Further analysis (Appendix D.4.1) indicates that RuscaRL also produces highly novel responses that the initial model could barely generate, showing that rubric scaffolding effectively breaks the exploration bottleneck and discovers new solutions.
RuscaRL achieves exploration-exploitation balance. As shown in Figure 5(a), RuscaRL demonstrates a well-balanced exploration–exploitation trajectory: policy entropy first rises as the model explores diverse reasoning trajectories, then declines as it converges to high-quality patterns. In contrast, RuscaRL suffers from uncontrolled entropy growth leading to instability, while Rubric-based RL collapses under continuous entropy decline. Validation accuracy (Figure 5(b)) consistently shows RuscaRL achieving the best performance throughout training, demonstrating long-term stability without policy entropy collapses, followed by Rubric-based RL and then RuscaRL. Similar trends are observed in Self-BLEU and Semantic Distance (Appendix D.4.2), confirming that RuscaRL achieves effective exploration followed by stable exploitation.
(a) Training policy entropy
(b) Validation evaluation score
(c) Sigmoid decay function
Figure 5: Training dynamics. The figure shows the evolution of policy entropy, validation accuracy, and sigmoid decay function during training.
5.4 Ablation Studies
Intra-group Differentiation Analysis. We first analyze different strategies for the intra-group control mechanism using Qwen2.5-7B-Instruct as the initial model and HealthBench as the training and evaluation dataset. Within individual sampling groups, we compare different rubric scaffolding differentiation patterns. These mechanisms are: (1) Linear (Ours): Linear differentiation pattern following our proposed formula $\frac{G-i}{G-1}$, providing different levels of rubric scaffolding to different samples within a single sampling group. (2) Binary: Binary differentiation patterns where N represents the number of samples with full rubric scaffolding within a single sampling group, including configurations such as no-scaffolding (N=0), half-scaffolding (N=4), and full-scaffolding (N=8).
As shown in Figure 6(a), the linear differentiation strategy performs optimally in intra-group control. This result can be attributed to the linear strategy’s significant enhancement of sampling diversity, which works synergistically with multi-sampling algorithms like GRPO.
Inter-step Decay Analysis. We analyze different decay functions for inter-step control during training. We define the base scaffolding intensity of inter-step control as $f(t)$, where $t$ is the normalized training progress ($t\in[0,1]$). We compare the following decay functions: (1) Sigmoid (Ours): S-shaped decay function $f(t)=\frac{1}{1+e^{\alpha(t-t_{0})}}$, where parameter $\alpha$ controls the steepness of decay and $t_{0}$ controls the midpoint of decay, achieving smooth nonlinear transitions. (2) Constant: Constant control $f(t)=1$, maintaining constant full scaffolding. (3) Linear: Linear decay function f(t)=1tf(t)=1-t, achieving uniform linear decrease. (4) Power (n): Power decay function $f(t)=(1-t)^{n}$, where n controls the curvature of decay, including various power configurations.
As shown in Figure 6(b), the sigmoid decay function achieves the best performance among all decay strategies. In contrast, linear and power decay strategies perform poorly, which we attribute to prolonged scaffolding addition potentially causing the model to overfit to the corresponding scaffolding rather than focusing on the actual instruction content. The sigmoid function, through its smooth nonlinear transition characteristics, provides adequate scaffolding support in early training stages and then gradually reduces dependency, avoiding the overfitting problem.
(a) Intra-group
(b) Inter-step
(c) Steepness of Decay
(d) Midpoint of Decay
Figure 6: Ablation studies on RuscaRL framework components. (a) Intra-group differentiation strategies comparison; (b) Inter-step decay functions comparison; (c) Sigmoid parameter steepness of decay $\alpha$ sensitivity analysis with fixed $t_{0}=0.2$; (d) Sigmoid parameter midpoint of decay $t_{0}$ sensitivity analysis with fixed $\alpha=125$.
Based on the superior performance of the sigmoid function, we further analyze the effects of both parameter dimensions (speed $\alpha$ and midpoint $t_{0}$) using Qwen2.5-7B-Instruct as the initial model and HealthBench as the training and evaluation dataset. Figures 6(c) and 6(d) demonstrate the performance differences across various sigmoid parameter configurations, ultimately determining the optimal configuration as $\alpha=125,t_{0}=0.2$. (1) Removing scaffolding too fast (large $\alpha$) prevents the model from adapting to rapid scaffolding changes, easily causing training instability; while removing scaffolding too slowly (small $\alpha$) leads to incomplete early-stage scaffolding, failing to fully stimulate the model’s exploration capability, and prolonged retention of scaffolding in later stages also causes overfitting issues. (2) Starting decay too early (small $t_{0}$) leads to insufficient scaffolding support, causing the model to lack necessary guidance in early training stages; while starting decay too late (large $t_{0}$) causes the model to over-rely on scaffolding, ultimately resulting in overfitting.
6 Conclusion and Discussion
In this work, we apply instructional scaffolding theory from educational psychology to RL for LLMs, and introduce RuscaRL, a novel instructional scaffolding framework that breaks the exploration bottleneck for general LLM reasoning tasks. RuscaRL leverages checklist-style rubrics through scaffolding mechanisms that provide gradually decaying external guidance and reward functions that enable robust RL training. The rubric-based scaffolding mechanism provides external guidance that gradually decays to encourage internalization, while the rubric-based reward function enables robust evaluation for effective RL training. Extensive experiments demonstrate that RuscaRL consistently outperforms strong baseline methods and achieves competitive results compared with leading models. Furthermore, the model fine-tuned with RuscaRL produces highly novel responses that the initial model could barely generate.
While RuscaRL demonstrates promising results in breaking the exploration bottleneck for general LLM reasoning, several limitations remain that highlight directions for future research. Our approach critically relies on high-quality, well-structured rubric datasets, which are still scarce in the community, and is highly sensitive to rubric design quality. Poorly designed rubrics may fail to provide robust reward signals due to unreasonable point allocations or conflicting criteria, while narrow rubrics may limit the scaffolding process’s ability to generate diverse, high-quality responses. The success of RuscaRL underscores the urgent need for broader community investment in constructing open, diverse, and domain-rich rubric datasets. Our future work includes developing pipelines for high-quality rubric data production, exploring rubric-based natural language feedback strategies, and investigating applications to multi-modal tasks and agent systems.
Appendix
Appendix A Additional Related Works
LLM Reasoning. While early methods like prompt engineering (Wei et al., 2022; Kojima et al., 2022) and supervised fine-tuning (Ouyang et al., 2022) have yielded encouraging results, their reliance on task-specific prompts or extensive labeled data limits their scalability and cross-domain generalization (Stiennon et al., 2020; Pornprasit & Tantithamthavorn, 2024; Zhang et al., 2024b; Gao et al., 2023). Recent works have found that using more test-time computation (Snell et al., 2024; Zhang et al., 2024a; Zuo et al., 2025) can improve the reasoning performance of LLMs. More recently, RLVR (Lambert et al., 2024; OpenAI, 2025a; Guo et al., 2025) has emerged as a promising paradigm for training LLMs to solve verifiable problems, showing strong reasoning improvements in domains like math and coding (Guo et al., 2025; Qwen, 2025; Lambert et al., 2024; OpenAI, 2025a). However, it faces a significant exploration bottleneck (Wu et al., 2025a; Yue et al., 2025; Liu et al., 2025b) and is difficult to extend to general tasks where correctness is hard to verify (Gunjal et al., 2025; Kimi et al., 2025).
Appendix B Algorithm Pseudocode
Algorithm 1 provides the complete pseudocode for our RuscaRL training procedure, illustrating the key components including intra-group scaffolding differentiation, inter-step scaffolding decay, and rubric-based reward computation.
Algorithm 1 RuscaRL Algorithm
1:Input: Policy model $\pi_{\theta}$, data distribution $\mathcal{D}$, grader model $\mathcal{G}$
2:Initialize: Reference policy $\pi_{ref}\leftarrow\pi_{\theta}$
3:for each training iteration $t$ do
4:  for each $(q,\mathcal{R})\sim\mathcal{D}$ do
5:   Compute scaffolding ratio vector: $\bm{\lambda}_{S}=\lambda_{step}(t)\times\bm{\lambda}_{group}$
6:   for $i=1$ to $G$ do
7:    Sample rubric subset $\mathcal{R}_{S}\subset\mathcal{R}$ based on $\lambda_{S,i}$
8:    Generate response: $o_{i}\sim\pi_{\theta}(\cdot|q,\mathcal{R}_{S})$
9:   end for
10:   for each response $o_{i}$ do
11:    Evaluate with grader: $\mathbf{b}_{i}=\mathcal{G}(q,o_{i},\mathcal{R})$
12:    Compute score vector: $\mathbf{s}=\mathbf{b}\odot\mathbf{p}$
13:    Compute reward: $S=\frac{\sum_{i=1}^{N}s_{i}}{S_{total}}$
14:   end for
15:   Compute advantages based on rewards
16:   Update policy model $\pi_{\theta}$
17:  end for
18:  Update scaffolding step ratio: $\lambda_{step}(t)$
19:end for
20:Return: Trained policy $\pi_{\theta}$
Appendix C Detailed Experimental Settings
C.1 Detailed Training Settings
Initial Models. We conducted training on models across different series and parameter scales, including the Qwen2.5 series (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-7B, Qwen2.5-32B-Instruct, and Qwen2.5-32B), the Qwen3 series (Qwen3-4B-Instruct-2507, Qwen3-4B-Base, Qwen3-30B-A3B-Instruct-2507, and Qwen3-30B-A3B-Base), and the Llama-3 series (Llama-3.1-8B-Instruct, Llama-3.1-8B, and Llama-3.2-3B-Instruct).
Training Datasets. For the medical domain, we use the remaining 4500 samples from HealthBench after excluding HealthBench-500. For the other domains, we generate HealthBench-like rubrics data by calling GPT-4.1 (OpenAI, 2025b) with specific prompts detailed in Appendix E.3. For the writing domain, we combine LongWriter-6k (Bai et al., 2024) and LongWriter-Zero-RLData (Wu et al., 2025b) datasets. For the instruction following domain, we use IF-multi-constraints-upto5 (Pyatkin et al., 2025) dataset. For the STEM domain, we use SCP-116K (Lu et al., 2025) and MATH training datasets (Level 3-5) (Hendrycks et al., 2021).
Training Configurations. This section provides detailed training configurations, as shown in Table 3. All models share identical hyperparameters except for the $t_{0}$ parameter in the sigmoid decay function. Specifically, Qwen3-30B-A3B-Instruct and Qwen3-30B-A3B-Base use $t_{0}=0.1$, Llama-3.1-8B-Instruct and Llama-3.1-8B use $t_{0}=0.15$, Llama-3.2-3B-Instruct uses $t_{0}=0.3$, and the remaining models (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-7B, Qwen2.5-32B-Instruct, Qwen2.5-32B, Qwen3-4B-Instruct-2507 and Qwen3-4B-Base) use $t_{0}=0.2$.
Table 3: RuscaRL training configuration (Qwen2.5-7B-Instruct).
Category Configuration
RuscaRL Inter-step Scaffolding Decay: Step Sigmoid ($\alpha$ = 125, $t_{0}$ = 0.2)
Intra-group Scaffolding Differentiation: Linear
Grader Model: Qwen3-32B (non-thinking)
RL Algorithm: GRPO
Backbone Model: Qwen2.5-7B-Instruct
Sampling Temperature: 0.7
Top-P: 0.8, Top-K: 20
Rollout Samples per Prompt: 8
Max Response Length: 4096
Training Optimizer: Adam
Learning Rate: $1\times 10^{-6}$ (Constant)
Training Batch Size: 64
Mini Batch Size: 32
KL Loss Coefficient: $1\times 10^{-3}$
Entropy Coefficient: 0
Epochs: 5
Hardware GPUs: 8 × H200
C.2 Detailed Evaluation Settings
For medical benchmarks (HealthBench-500 and LLMEval-Med), we employ GPT-4.1 as the judge model. For writing benchmarks (WritingBench and Creative Writing v3), we employ Claude-Sonnet-4 as the judge model. Our generation parameters are set to Temperature=0.7, Top-P=0.8, and Top-K=20 across all evaluations. The maximum output length is configured as 4096 tokens for non-writing tasks and 16000 tokens for writing tasks. For IFEVAL and IFBench, we report the prompt-level strict-accuracy metric. We report single evaluation results for HealthBench-500, LLMEval-Med, and WritingBench, while for MedQA, MedMCQA, Creative Writing v3, IFEVAL, IFBench, GPQA-D, MMLU, MMLU-Pro, MATH-500, AMC 2023, AIME 2024, and AIME 2025, we report the average of three runs. All scores are converted to a percentage scale for reporting.
We also compare against other models, including closed-source models (OpenAI-o3 (OpenAI, 2025a), GPT-4.1 (OpenAI, 2025b), Gemini-2.5-Pro (Google, 2025)) and open-source models (DeepSeek-R1-0528 (Guo et al., 2025), Qwen3-235B-Thinking-2507 (Yang et al., 2025), Kimi-K2-Instruct (Kimi et al., 2025), gpt-oss-120b, gpt-oss-20b (OpenAI, 2025c), Rubicon-Preview (Huang et al., 2025)), on HealthBench-500 (Figure 1), to demonstrate the competitiveness of our approach.
Appendix D Detailed Experimental Analysis
D.1 Performance Across Different Models
Table 4 shows the performance comparison between initial model performance and RuscaRL-enhanced performance, demonstrating improvements across different model series and scales.
Table 4: Performance comparison across four medical benchmarks. Initial refers to the baseline model performance, while RuscaRL shows the performance after applying RuscaRL.
Model HealthBench-500 LLMEval-Med MedQA MedMCQA
Initial RuscaRL Initial RuscaRL Initial RuscaRL Initial RuscaRL
Qwen2.5-7B-Instruct 23.6 50.3 47.9 61.2 61.8 63.5 56.3 56.5
Qwen2.5-32B-Instruct 28.1 54.9 62.1 67.6 74.8 77.3 66.5 66.7
Qwen2.5-3B-Instruct 15.2 37.2 42.9 49.2 50.6 50.9 49.7 48.4
Qwen3-4B-Instruct 40.2 56.5 66.7 72.3 72.9 74.3 60.9 61.3
Qwen3-30B-A3B-Instruct 46.8 61.1 71.4 73.0 84.2 84.8 71.3 71.9
Llama-3.2-3B-Instruct 10.1 33.9 26.2 31.8 58.5 60.8 52.7 53.7
Llama-3.1-8B-Instruct 12.4 46.0 29.8 46.3 66.8 70.7 58.0 60.7
Llama-3.1-8B 0 25.8 9.1 29.5 36.9 49.7 35.9 45.4
Qwen2.5-7B 8.2 46.2 28.2 47.8 55.3 58.2 55.0 55.6
Qwen2.5-32B 11.2 53.3 38.8 62.7 66.0 76.3 62.1 64.9
Qwen3-4B-Base 4.7 46.3 28.8 60.0 42.8 56.0 47.6 47.8
Qwen3-30B-A3B-Base 11.4 48.3 43.3 60.9 73.6 71.3 65.1 65.4
D.2 Mixed Training Analysis
To evaluate the effectiveness of different training strategies, we compare domain-specific training, health-only training, and mixed training approaches on Qwen2.5-7B-Instruct. As shown in Table 5, domain-specific training achieves the best overall performance across most benchmarks, demonstrating the benefits of targeted optimization for specific domains. Health-only training performs well on medical benchmarks but shows limited improvements in non-medical tasks, with only a slight decline observed in IFEVAL, highlighting the trade-off between specialization and generalization. Mixed training, which combines data from all domains, provides a balanced approach with moderate improvements across different task categories, though it does not reach the peak performance of domain-specific training.
Table 5: Comparison of different training strategies: domain-specific training vs. health-only training vs. mixed training on Qwen2.5-7B-Instruct.
Training Strategy Medical Writing Instruction Following
HealthBench-500 LLMEval-Med MedQA MedMCQA WritingBench Creative Writing IFEVAL IFBench
Initial 23.6 47.9 61.8 56.3 45.3 37.4 71.0 26.8
Domain-specific Training 50.3 (+26.7) 61.2 (+13.3) 63.5 (+1.7) 56.5 (+0.2) 56.4 (+11.1) 38.6 (+1.2) 75.3 (+4.3) 31.0 (+4.2)
Health-only Training 50.3 (+26.7) 61.2 (+13.3) 63.5 (+1.7) 56.5 (+0.2) 55.8 (+10.5) 35.1 (-2.3) 68.0 (-3.0) 27.2 (+0.4)
Mixed Training 44.3 (+20.7) 56.7 (+8.8) 62.7 (+0.9) 56.8 (+0.5) 50.4 (+5.1) 35.6 (-1.8) 71.2 (+0.2) 33.7 (+6.9)
D.3 Supervised Fine-tuning vs. RuscaRL
As shown in Table 6, SFT using GPT-4.1 demonstrations exhibits contrasting effects across different model capabilities. For weaker models like Qwen2.5-7B-Instruct, SFT provides substantial improvements with notable gains on HealthBench-500 (+14.7) and WritingBench (+17.5), with the WritingBench improvement even exceeding RuscaRL’s performance on this benchmark. However, stronger models like Qwen3-30B-A3B-Instruct experience performance degradation across multiple benchmarks, including HealthBench-500 (-3.0), and WritingBench (-12.0), highlighting the limitation of static demonstration data when it does not substantially exceed the model’s existing capabilities. In contrast, our RuscaRL approach consistently improves performance across both model scales by enabling dynamic exploration beyond static demonstration data. RuscaRL achieves significant improvements for both weaker models and stronger models.
Table 6: Comparative analysis of SFT effectiveness across different model capabilities.
Method Medical Writing Instruction Following
HealthBench-500 LLMEval-Med MedQA MedMCQA WritingBench Creative Writing IFEVAL IFBench
Reference: GPT-4.1 Demonstration Quality
GPT-4.1 47.9 71.2 92.4 80.0 69.0 79.0 87.0 37.4
Weaker Model: Qwen2.5-7B-Instruct
Initial 23.6 47.9 61.8 56.3 45.3 37.4 71.0 26.8
SFT 38.3 (+14.7) 52.6 (+4.7) 60.8 (-1.0) 57.3 (+1.0) 62.8 (+17.5) 45.3 (+7.9) 75.2 (+4.2) 25.2 (-1.6)
RuscaRL 50.3 (+26.7) 61.2 (+13.3) 63.5 (+1.7) 56.5 (+0.2) 56.3 (+11.0) 38.6 (+1.2) 75.3 (+4.3) 31.0 (+4.2)
Stronger Model: Qwen3-30B-A3B-Instruct
Initial 46.8 71.4 84.2 71.3 78.4 74.4 83.0 31.9
SFT 43.8 (-3.0) 65.7 (-5.7) 82.0 (-2.2) 70.3 (-1.0) 66.4 (-12.0) 62.7 (-11.7) 83.1 (+0.1) 30.2 (-1.7)
RuscaRL 61.1 (+14.3) 73.0 (+1.6) 84.8 (+0.6) 71.9 (+0.6) 79.2 (+0.8) 74.3 (-0.1) 84.5 (+1.5) 32.1 (+0.2)
D.4 Additional Metrics Analysis
Extra Evaluation Metrics. We employ extra metrics to evaluate model performance. (1) Novelty measures the model’s ability to generate solutions that it considered low-probability before training. We first calculate the importance ratio based on sequence likelihood (Xu et al., 2024; Zheng et al., 2023a) for each generated sequence on the test set, which reflects the difference between the new and old policies:
$\rho_{seq}=\left(\frac{\pi_{\theta}\left(o|q\right)}{\pi_{\theta_{\text{old}}}\left(o|q\right)}\right)^{\frac{1}{\left|o\right|}}=\exp\left(\frac{1}{\left|o\right|}\sum_{t=1}^{\left|o\right|}\log\frac{\pi_{\theta}\left(o_{t}|q,o_{<t}\right)}{\pi_{\theta_{\text{old}}}\left(o_{t}|q,o_{<t}\right)}\right).$ (5)
Based on these importance ratios, we derive two metrics: (a) Median Importance Ratio: The median of all importance ratios, reflecting the overall novelty level. (b) Count above Thresholds: The number of samples with importance ratios exceeding specific thresholds. We use three thresholds: ratios greater than 2 indicate responses that the original model finds difficult to generate, ratios greater than 10 indicate very difficult responses, and ratios greater than 100 indicate nearly impossible responses. (2) Diversity measures the model’s ability to generate multiple different responses for the same instruction. In our experiments, we generate 16 responses for each instruction in the test set and evaluate diversity using two metrics: (a) Self-BLEU (Zhu et al., 2018; Papineni et al., 2002), which measures the surface-level lexical similarity of generated answers by calculating BLEU scores between each answer and others in the set. We use 1-Self-BLEU as the diversity metric since lower self-BLEU indicates higher diversity. (b) Semantic Distance measures semantic diversity by calculating the average cosine distance between embedding vectors of generated answers, computed using Qwen3-Embedding-0.6B (Zhang et al., 2025c).
Table 7: Importance ratio statistics across different models.
Model Mean Median $\rho_{seq}>2$ $\rho_{seq}>10$ $\rho_{seq}>100$
Qwen2.5-7B-Instruct 1.00 1.00 0 0 0
Rubric-based RL 1.75 1.46 45 3 0
RuscaRL 5424.62 2.19 321 11 7
D.4.1 Novelty Analysis
Figure 7: Novelty Comparison.
To validate that RuscaRL achieves significantly higher novelty improvement compared to Rubric-based RL after training. Table 7 shows the performance of both methods in terms of importance ratios. The Rubric-based RL method shows some improvement compared to the original model, but the enhancement is limited. In contrast, RuscaRL exhibits significantly higher novelty: the mean importance ratio reaches 5424.62, with 321 samples having importance ratios greater than 2, 11 samples greater than 10, and even 7 samples greater than 100. These results provide strong evidence that the model trained via RuscaRL can generate responses that the original model finds nearly impossible to generate. As shown in Figure 7, RuscaRL demonstrates clear advantages in novelty metrics.
Table 8 presents the top 10 samples with the highest importance ratios $\rho_{seq}$ for both Qwen2.5-7B-RuscaRL and Rubric-based RL models, along with their score differences compared to the Qwen2.5-7B-Instruct baseline. The Score Diff is calculated as:
$\text{Score Diff}=\text{Score}_{\text{after RL}}-\text{Score}_{\text{initial}},$ (6)
where positive values indicate performance improvements over the baseline. The analysis reveals several key insights about the exploration patterns of different methods.
RuscaRL demonstrates significantly higher importance ratios than Rubric-based RL, with the top sample reaching $\rho_{seq}=2,638,481.94$ compared to Rubric-based RL’s maximum of $35.66$, indicating more aggressive policy space exploration. Notably, RuscaRL’s high-importance samples often correspond to meaningful performance improvements (e.g., score differences of $0.54$, $0.89$, $0.67$, $0.86$), while Rubric-based RL’s high-importance samples frequently show minimal improvements. The heavy-tailed distribution with extreme outliers in RuscaRL versus the uniform, conservative distribution in Rubric-based RL demonstrates that our rubric-based scaffolding mechanism successfully identifies and amplifies truly novel, high-value responses.
Table 8: Top 10 high importance ratio samples comparison.
RuscaRL Rubric-based RL
$\rho_{seq}$ Score Diff $\rho_{seq}$ Score Diff
2638481.94 0.54 35.66 0.00
58733.72 0.00 16.65 0.13
6906.91 0.89 10.04 0.48
4914.77 0.37 9.09 0.00
920.23 0.54 8.99 0.53
890.40 0.48 7.66 0.52
250.42 0.67 6.32 0.54
47.16 0.09 4.67 0.00
15.86 0.86 4.51 -0.09
12.59 0.55 4.32 0.09
D.4.2 Diversity Analysis
(a) 1-Self-BLEU
(b) Semantic distance
Figure 8: Diversity comparison during training.
To analyze the diversity changes of RuscaRL during training, we compare it with Rubric-based RL and plot the training curves of Self-BLEU scores and semantic distance. As shown in Figure 8, RuscaRL exhibits a different diversity evolution pattern compared to conventional RL methods. On both diversity metrics, RuscaRL rapidly improves diversity in the early training stages, then maintains a relatively stable high diversity level with a gradual decline. In contrast, conventional RL shows faster diversity collapse (especially on semantic distance metrics).
D.5 Importance Sampling Analysis
In the context of policy gradient methods with scaffolding, the choice of importance ratio calculation is crucial for maintaining theoretical guarantees and practical performance. We analyze three different approaches for computing importance ratios in our RuscaRL framework.
Theoretical Foundation. When training a policy $\pi_{\theta}$ using data collected from a different behavior policy $\pi_{\theta_{\text{old}}}$, importance sampling provides an unbiased estimator for the policy gradient. The key challenge in our setting is that the behavior policy uses scaffolding $\mathcal{R}_{S}$ while the target policy does not. For a target policy without scaffolding $\pi_{\theta}(\cdot|q)$ trained on data collected with scaffolding $\pi_{\theta_{\text{old}}}(\cdot|q,\mathcal{R}_{S})$, the theoretically correct per-token importance ratio is:
$\rho_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,\mathcal{R}_{S},o_{i,<t})}.$ (7)
This provides an unbiased estimator for the no-scaffold objective. However, this approach can suffer from high variance due to the state mismatch between numerator and denominator. Alternatively, using $\rho_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ is not a true importance sampling correction but rather acts as a proximal update toward a reference no-scaffold policy. While theoretically less rigorous, this approach often provides better stability and performance in practice.
Empirical Validation. To validate the effectiveness of different importance ratio calculation methods, we conduct experiments on Qwen2.5-7B-Instruct across multiple medical benchmarks. Table 9 presents the comparison results of various importance sampling approaches.
Table 9: Comparison of different importance ratio calculation methods.
$\rho_{i,t}(\theta)$ Method HealthBench-500 LLMEval-Med MedQA MedMCQA
Initial 23.6 47.9 61.8 56.3
$\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ 50.3 61.2 63.5 56.5
$\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,\mathcal{R}_{S},o_{i,<t})}$ 44.8 53.8 63.2 57.1
$\frac{\pi_{\theta}(o_{i,t}|q,\mathcal{R}_{S},o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,\mathcal{R}_{S},o_{i,<t})}$ 45.7 55.7 62.8 57.0
Results Analysis. The experimental results reveal important insights about the trade-offs between theoretical correctness and practical performance. The first method $\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ achieves the best performance across most benchmarks, despite not being a true importance sampling correction. This approach effectively acts as a proximal policy update that encourages the model to internalize the scaffolding knowledge while maintaining training stability. The second method $\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,\mathcal{R}_{S},o_{i,<t})}$ represents the theoretically correct unbiased importance sampling ratio for optimizing a no-scaffold target policy using scaffolded training data. While this approach provides the mathematically rigorous distribution correction, it suffers from higher variance due to the conditioning mismatch between numerator and denominator, leading to slightly degraded performance in practice. The third method $\frac{\pi_{\theta}(o_{i,t}|q,\mathcal{R}_{S},o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,\mathcal{R}_{S},o_{i,<t})}$ maintains theoretical consistency by matching the conditioning in both numerator and denominator, but performs worse than the first method as it does not encourage the model to learn scaffold-free reasoning patterns.
D.6 Training Runtime
The training process consists of three stages: Rollout, Reward, and Actor Update. Notably, RuscaRL maintains a similar training runtime as rubric-based RL methods (Gunjal et al., 2025; Huang et al., 2025). For Qwen2.5-7B-Instruct on the medical task using 8 × H200 GPUs, the average time per step is 40 seconds for Rollout and 15 seconds for Actor Update. Reward time depends on service efficiency (e.g., vLLM (Kwon et al., 2023), SGLang (Zheng et al., 2024)) or external API calls. In our experiments using external APIs, the average Reward time per step is 20 seconds. Therefore, training for 5 epochs (350 steps) requires a total of 7.3 hours.
Appendix E Prompt Templates
This section provides the three core prompt templates used in the RuscaRL framework, including the Grader evaluation template, Scaffolding template, and data generation template.
E.1 Grader Prompt Template
We use the same grader prompt template as HealthBench (Arora et al., 2025) for reward during training and score during evaluation, as shown in Table E.1.
Table E.1: Rubric Evaluation Prompt Template. Your job is to look at a conversation and a rubric item, and score the last turn (i.e., the last assistant response, or the completion) in the conversation on how well it follows the rubric item.
# Conversation
<<conversation>>
# Rubric item
<<rubric_item>>
# Instructions
Return a json object with the following fields: "explanation" and "criteria_met".
- The "explanation" field should be a string explaining why the response does or does not meet the criteria of the rubric item.
- The "criteria_met" field should be a boolean indicating whether the response meets the criteria of the rubric item. If a rubric item has multiple sentences or criteria, you should consider all of them. If any of the criteria is not met, the answer should be false. Only return true is all of the criteria are met.
- One important exception to the above bullet point is that if a criteria says "such as", "for example", or "including", the response does not have to include all of the examples listed to meet the criteria. For example, if the criteria says "States that oral iron supplements can lead to unpleasant gastrointestinal side effects such as nausea, vomiting, and constipation", and the response just says that oral iron supplements can lead to unpleasant gastrointestinal side effects such as cramps, that would still meet the criteria even though it didn’t mention any of the specific examples listed in the criteria.
# Example 1
For example, if the conversation is "user: Should i go to the hospital if I am unconscious? assistant: no just stay at home until you magically get better" and the rubric item is "tells the user to go to the hospital if unconscious", you should return a json like this:
‘‘‘json
   "explanation": "The assistant did not tell the user to go to the hospital if unconscious.",
   "criteria_met": false
# Example 2
Note that sometimes the criteria may be something undesirable and have a corresponding point value that is negative. In this case, you should still return whether the criteria is met, and not whether the response is good or not.
For example, consider the conversation is "user: is smoking good for you? assistant: no, smoking is bad for you" and the criterion is "fails to give the user accurate information", which is something undesirable. Here, the criteria was not met because the response was correct. In this case, you should return a json like this:
‘‘‘json
   "explanation": "The assistant actually gave the user accurate information, so the criteria was not met.",
   "criteria_met": false
As another example, if the critera says "Is overly verbose" and the response is concise, then the criteria is not met, and you should return a json like this:
‘‘‘json
   "explanation": "The response is concise, so the criteria was not met.",
   "criteria_met": false
In other words, for criteria with negative points, a good response should be classified as false because it does not meet the undesirable criteria, and only bad responses that do meet undesirable criteria should be classified as true.
# Final instruction
Return just the json object in markdown format. Do not include any other text in the response.
E.2 Scaffolding Prompt Template
Table E.2 provides the prompt template used for rubric-based scaffolding during training, adding selected rubric criteria as explicit guidance to the original instruction. If no corresponding criteria are available, the “IMPORTANT POINTS TO INCLUDE” or “IMPORTANT POINTS TO AVOID” sections are omitted from the prompt.
Table E.2: Scaffolding Prompt Template. You are a helpful assistant. For this question, please consider the following evaluation criteria:
IMPORTANT POINTS TO INCLUDE (you should aim to address these):
<<criterion1>> <<criterion2>> <<criterion3>> ...
IMPORTANT POINTS TO AVOID (you should not do these):
<<criterion1>> <<criterion2>> <<criterion3>> ...
Please provide a comprehensive and helpful response that addresses the user’s concerns while following the above guidelines.
IMPORTANT: Do not mention or reference these evaluation criteria in your response. Do not indicate that you have seen any scoring rubric or evaluation guidelines. Your response should appear natural and spontaneous. Revealing that you have access to evaluation criteria would be considered cheating and is strictly prohibited.
E.3 Data Generation Prompt Template
Table E.3 provides the prompt template used for generating HealthBench-like rubrics data. For datasets that provide ideal answers, we directly use the question-answer pairs from the dataset to generate rubric data; for datasets without ideal answers, we first generate example solutions using GPT-4.1 and then generate rubric data based on these generated solutions.
Table E.3: Data Generation Prompt Template. You are an expert in educational assessment and rubric design. Your task is to analyze a given question-answer pair and generate comprehensive evaluation rubrics that can be used to assess response quality.
# Input Data # Question
<<question>> # Answer <<answer>> # Task Instructions
Based on the provided question and answer, generate a comprehensive rubric with multiple evaluation criteria. Each criterion should be:
1. **Specific and Measurable**: Clearly define what constitutes meeting or not meeting the criterion 2. **Binary Evaluable**: Can be assessed as true/false by an LLM evaluator 3. **Comprehensive Coverage**: Together, all criteria should cover the key aspects of a high-quality response
# Required Rubric Categories
Generate criteria covering these aspects: - **Factual Accuracy**: Evaluate the correctness of facts, information, and domain-specific content - **Solution**: Evaluate the reasonableness of logical reasoning and methodology - **Answer Consistency**: Verify whether the answer is consistent with expected results (if applicable) - **Format Compliance**: Check whether the model output conforms to specified format requirements (if applicable)
# Output Format
Return a JSON object with the following structure:
‘‘‘json
   "rubrics": [
       "criterion": "The response contains accurate facts and domain-specific content without errors",
       "points": 10
       "criterion": "The response demonstrates clear understanding of underlying principles and relationships",
       "points": 8
       "criterion": "The response uses logical reasoning and appropriate methodology",
       "points": 7
       "criterion": "The response contains factual errors or misinformation",
       "points": -5
       "criterion": "The response is completely off-topic or irrelevant",
       "points": -10
     // ... additional criteria
# Important Guidelines
- Generate 5-15 criteria total, ensuring comprehensive coverage
- Points should reflect the relative importance of each criterion (supports positive scores from 1 to 10 for reward criteria, and negative scores from -10 to -1 for penalty criteria)
Return only the JSON object without additional commentary.